# Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling

### v1

@@ 4.2. Standards Identification & Management
Auditors also used tools to formulate the principles and norms to guide their investigations. This stage, Standards Identification & Management ($N=184$), includes tools for Goal Articulation (e.g., broad principles statements), Self-Assessment (more specific procedural assessment tools, such as Microsoft’s AI Fairness Checklist (madaio_co-designing_2020, 65)), Documentation (e.g., Model Cards (mitchell_model_2019, 68)), Regulatory Awareness (tools, often paid services, for discovering and monitoring relevant regulations), Methods Design (standards for audit methodology), and Participatory Standard-Setting (methods for developing standards in collaboration with affected groups, such as Microsoft’s Community Jury (cass_community_2022, 19)).
Standards Identification & Management comprised the largest collection of tools in our database. In particular, this stage includes the “principle statements” and similar documents ($N=75$) that many organizations—particularly nonprofits and government agencies (77.2% not-for-profit)—have created to articulate their broad goals for AI systems and AI procurement. Checklists and other forms of procedural self-assessment, intended for developers and researchers, were also common ($N=41$), while methods for participatory standard-setting ($N=6$) were comparatively rare. We also noted the format of each tool we found, annotating when the tool included software (either a software product or an API), a code repository, a data repository, a white paper/report, or a guidebook (Fig. C.6). While every other category has more tools involving code or data than without, only 8% of Standards Identification tools included code or data.
@@
Need for more standardized, context-specific, and holistic evaluation frameworks. Despite the large number of standards and evaluation frameworks surfaced in our landscape analysis, auditors still felt that evaluation frameworks needed refinement. Auditors emphasized the importance of standardized evaluation frameworks that provide clarity and consistency: “I think standardization is a big [concern]…” (P23) Many expressed a vision to streamline the auditing process by offering predefined structures and templates for assessment, which are essential for conducting audits effectively and facilitating communication.
Auditors wished that standards were both more context-specific and more holistic. Standards identification tools were particularly general in their applications, compared to other categories of tools—principle statements, checklists, and similar resources were usually intended for multiple kinds of audit targets (Fig. C.5). Some participants found these tools too vague:
@@
I think what I’ve seen is that companies and institutions… really, really struggle to understand, ‘What should we even do when it comes to auditing or evaluating the use of machine learning in our organizations?’ And while a template or a checklist is not the right answer, a lot of them don’t even know where to start… And so, (it helps) when you have a tool that…has some built in frameworks, and maybe there’s some big ones in there, so like maybe the NIST framework (tabassi_ai_2023, 99) is in there. (P8)
@@
Despite generality in application, though, some participants thought that assessments templates and checklists were focused too narrowly on fairness assessment. One civil society auditor said, “I would love some guidance on audits generally and tools that describe the non-fairness components of an audit…” (P6) Another described collating multiple standards themselves, wishing for a resource that covered criteria such as explainability, privacy, and transparency: “Currently we have to.. find open source tools and put them together ourselves, and you need to have expertise to know what to look for” (P21).
@@
Need for regulatory guidance. Despite some participants’ desire for standardized frameworks like NIST’s (tabassi_ai_2023, 99), multiple participants commented on the difficulty of harmonizing current or expected regulatory guidance with practical implementation. One civil society auditor said:
@@
There’s two different points…: What are we evaluating for? And the question of, even when we have some kind of legal or other benchmark in mind, what are the metrics, and what are the benchmarks and other ways in which to evaluate, technical and otherwise, which also remain quite unclear? We’re seeing the ready adoption of audit language into policies, so that just kind of makes us nervous… What are we auditing for? (P2)
@@
Several emphasized the necessity of regulatory entities being more forthcoming in defining best practices. Some sought the guidance of the regulatory bodies in the domains they operate to establish their own frameworks but often struggle to connect or translate industry expectations into meaningful standards for evaluation. Multiple participants felt there was a “culture or communications gap between the legal and compliance people on one side and the engineers on the other” (P12). As a result, some hoped for more collaborative approaches instead of command-and-control prescriptions:
Half the time we reach out, and there’s just no one to contact… and it’s just a black hole… [Another way forward is] not regulation, but it’s more like insurance liability and these sort of semi-regulatory structures that force you to engage… (P13)
@@
Implications. Standards for evaluating AI systems must be simultaneously holistic, context-specific, and compatible with practice. Some of the tools we found made advances in one or more of these dimension, but few accomplished all three. Microsoft’s AI Fairness checklist, for example, was co-designed with practitioners (madaio_co-designing_2020, 65) in an effort to be more compatible with practical challenges, but many of the other standards frameworks we found did not obviously consult practitioners—and even fewer involved affected stakeholders. Likewise, while NIST’s AI Risk Management Framework includes safety, security, reliability, transparency, explainability, and privacy, in addition to fairness, its guidance for specific evaluation techniques remains fairly broad (tabassi_ai_2023, 99).
@@
Future work could further explore how resources and tools (such as templates and documentation workflows) can generally make standards management easier and more relevant for practitioners while also involving affected stakeholders to ensure that standards remain inclusive and impactful. Research could continue to explore how regulatory standards could be translated into concrete metrics or other expectations for industry (wachter_bias_2021, 105), even as policymakers aim to provide more effective and feasible guidance (guha_ai_2023, 45). One participant at a large tech company, for example, preferred a “must, could, should” (P14) structure for regulatory guidance: a non-technical legal minimum (“must”) accompanied by more precise, technically feasible paths to compliance (“should”) and a set of ideal best practices for high performers and innovators (“could”). Research on the organizational practice of responsible AI (widder_its_2023, 109, 47) could help improve collaboration between standards-setting and practice.
@@ 4.3. Data Collection & Transparency Infrastructure
Gathering empirical evidence is a key step in AI auditing, but often posed a significant challenge—especially to independent auditors. When model operators were unwilling or unable to release relevant documentation and other evidence, auditors turned to two main classes of tools. First, tools for Transparency Infrastructure ($N=12$) are interfaces and databases hosted by model operators that allow controlled access to relevant data. This category includes tools for Structured or Application Programming Interface (API) Access (tools that allow auditors to interact with models and live systems in controlled ways, such as Google’s AI Test Kitchen (warkentin_join_2022, 107)), tools for Data Sharing (platforms or trusts for hosting models and related data, such as the Gig Economy Data Hub (noauthor_gig_2021, 41)), and tools for Secure & Private sharing (tools that help mitigate security and privacy concerns with sharing data, such as Airbnb’s Project Lighthouse (airbnb_new_2020, 3)).
@@
More commonly, though, tools for Data Collection ($N=42$), helped auditors gather information outside these controlled interfaces. These tools help auditors gather data about model behavior, including relevant information not routinely collected by model operators. This category includes tools for Field Data Collection, which collect data from real systems and real users—including tools for Data Donation (such as Mozilla’s YouTube Regrets project (mozilla_foundation_youtube_2021, 69)), Data Scraping (e.g., Tracking Exposed (agosti_tracking_2023, 1)), Interviews/Surveys, and Compelled Transparency (e.g. tools such as MuckRock that facilitate FOIA requests). We also found tools for Simulation (e.g. Meta’s Web Enabled Simulation platform for simulating interactions on Facebook (ahlgren_wes_2020, 2)) and Bot Deployment (tools used for sock puppet auditing (bandy_problematic_2021, 7), such as Selenium or Appium), both used to test systems in artificial or semi-artificial interactions.
Need for uncompromised data access. These tools have been used to help overcome one of the challenges most frequently mentioned by our participants: the difficulty of accessing data and other vital information required to conduct meaningful audits. While there are a few existing tools that provide external auditors with controlled access to models and data—especially for online platforms (Fig. C.5)—they require voluntary investment by model operators. One participant noted that key APIs used for auditing are becoming more costly and undependable: “There’s a direct impact on shutting off the access to information that affects people doing audits… we saw that with Reddit charging for their API and shutting down… Twitter charging astronomical, now, amounts for their API. It’s because everybody is scraping public Reddit and public Twitter to train large [AI] models…” (P24). They wished for “access to platform data, or like doing like a context under which people can do controlled experiments, using data that is provided by platforms directly” (P24). Another civil society auditor said, “I wish I had something to force people to give me their data… Part of the problem of auditing in general is that the only people who get let in are usually the people who are willing to say nice things about whatever the technology is being audited” (P6).
Auditors also said that corporate control over APIs created a lack of trust among investigators. Indeed, the tools for Transparency Infrastructure we found were disproportionately built by large, for-profit tool developers (75.0% built by for-profits), compared to tools for Data Collection (19.0% built by for-profits), which came more often from non-profits, government agencies and academics (Fig. C.7). In practice, participants reported that key details, such as data sampling methods, data provenance, model versioning, metrics, and design justifications, were often omitted or only partially disclosed.
@@
Currently, the vetting process for API use often requires the auditor to disclose their intent for the evaluation in advance, which may compromise the integrity of the study. “I’m very much concerned that what’s going to happen is.. [platforms have] given all this access, and there will actually be… more of a cover up than there is now… They have so much power in this conversation to just share whatever information they want” (P7). Instead, one participant’s ideal was an “inspectability API” (P7), a required, standardized interface to allow researchers to interact with online platforms, including the ability to test different profiles, geographies, and other variables needed to evaluate disparate treatment, misinformation, and other algorithmic harms. Similarly, an auditor at a startup wished for centralized data archive available to the public:
@@
Challenges with independent data collection. Rather than rely on the model operator to provide access, auditors—especially external auditors—often turned to tools for Data Collection to obtain evidence themselves, sometimes developing and sharing their own tools and processes. Unlike tools for Transparency Infrastructure, which were mostly proprietary (25.0% open source), tools for Data Collection we found were much more likely to be available under an open source license (73.8% open source; see Figure 3). Tools for Data Collection also included the most popular Github repositories in our survey (Table 4)—for example, auditors used Selenium (noauthor_selenium_nodate, 87), a popular collection of open source tools for browser automation, to simulate user profiles while scraping data (known as a “sock puppet” audit (bandy_problematic_2021, 7)).
(a) Number of open source tools in each taxonomy stage.
@@
(a) Number of open source tools in each taxonomy stage.
@@
Figure 3. Tool licensing by taxonomy stage (left) and by organization type (right).
While this approach gave auditors more freedom, it could also take more effort. Some of the tools we found—such as the Markup’s Citizen Browser (the_markup_citizen_2022, 100), a data donation platform—were built from scratch to collect specific kinds of data for auditing. Existing tools for data scraping were helpful but often required extensive adaptation.
We almost always have to build custom scrapers to collect data… There’s some templates right out there for these scrapers, and then you usually have to customize them. And then there’s a huge amount of like work to keep them alive… They break all the time with all these edge cases. And so they’re really a pain. I don’t really know that there’s a way to solve that. (P7)
Despite these difficulties, external auditors—particularly the journalists in our sample—still saw advantages in independent data collection. One journalist noted:
@@
Challenges with understanding and processing data. Even when they had access to the data they wanted, auditors also struggled with the basic challenges involved with managing and analyzing data. Participants noted that this stage of auditing often necessitates more manual labor than any other. A government auditor said, “90% of the work is figuring out what different tables are, and what different columns are, and working out whether it makes sense to join certain things. And then figuring out some meaningful metrics that we can draw from that data…” (P25). In addition to typical data science tasks involved with processing large datasets, participants also reported spending lots of time reviewing documentation and requesting additional information about audit targets.
Our survey did not surface any tools designed to assist with data management or data processing for AI auditing, nor did the audit reports we read or our participants mention any specific tools used for this purpose. However, multiple participants wished for tools to help with these tasks. One auditor expressed that infrastructure may be most useful for more time-consuming steps in the auditing process, such as data collection and cleaning, stating, “most of the value of data infrastructure is literally cleaning data” (P3). One auditor hoped for innovation in data quality management: “There are custom scrapers, a lot of human data quality work, and one thing that I have really wanted to do and never been able to do is try to figure out ways to get that data quality work done in more interesting ways” (P7). Participants repeatedly urged the development of standardized tooling and infrastructure to facilitate data collection and cleaning, in order to streamline the auditing process.
@@
Risk of retaliation. Some external auditors also worried that external data collection tools—particularly tools for data scraping or data donation—may violate terms of service set by platforms and result in legal liability or retaliation. In 2021, for example, Facebook disabled the accounts of a group of researchers auditing its advertising algorithm using a data donation tool called Ad Observer (edelson_we_2021, 34). Facebook initially insinuated that disabling the researchers’ accounts was required by an Federal Trade Commission (FTC) consent decree related to privacy, backtracking only after the FTC called the statement inaccurate (levine_letter_2021, 60).
In particular, auditors expressed concern about the broad and ambiguous language in existing laws and regulations—particularly the Computer Fraud and Abuse Act (CFAA), but also the General Data Protection Regulation (GDPR) and other privacy laws—that could expose them to legal risk. Under the CFAA, for example, an auditor who violates a platforms’ terms of service may be held criminally liable. As it stands, the CFAA does not include a list of clear exemptions for auditors doing research for the public interest. And it is difficult for auditors to determine what methods—such as public data scraping or even data donation—may be deemed “unauthorized” and criminal under the CFAA unless the audited organization grants explicit authorization. In Sandvig v. Barr (2019) the ACLU sued to allow researchers to set up false accounts (sock puppets) to audit computer algorithms (noauthor_sandvig_2020, 85), arguing that this provision exerts a chilling effect on auditing research (american_civil_liberties_union_sandvig_2019, 4). Indeed, Facebook used exactly this term of art—“unauthorized”—several times in its explanation for shutting the Ad Observer project accounts (clark_research_2021, 22). As one civil society auditor said, “Even if the legal risks have not been acted upon as much, there’s cases everyone points to in terms of attacks against researchers. It’s a matter of time before it ramps up. As soon as our work becomes threatening enough, that’s when it all really starts” (P24).
As a result, auditors engaged in external data collection had to take great steps to guard against legal liability and retaliation. One journalist said, “[The CFAA] is an incredible legal hangover for the type of work I do… how much lawyering I need to even get one tool off the ground is insane” (P7). Auditors also expressed hesitation about reforms that give platforms more control over what data is released. The same journalist continued,
@@
Implications. Given that data access and integrity is critical for auditing and for accountability, future research could explore tools and processes for not only facilitating access to data—especially independently, through scraping or simulation—but also for ensuring data quality and integrity. Participants specifically wished for more tooling for data donation and user-driven auditing—a nascent area of research in human-computer interaction (lam_sociotechnical_2023, 56, 32, 29). Auditors also faced challenges common to data work in general, and research on practices surrounding data quality and data integrity may be applied specifically to discrimination testing and AI auditing methods.
Other challenges were more particular to auditing work. Future work could explore, for example how auditors request information and interact with model operators. Transparency Infrastructure in particular is a nascent area of tooling that may become more common to auditing as AI regulation develops and as barriers to external data collection mount; independent research may help guide these tools into more trustworthy mechanisms for disclosure, even as policymakers can ensure platform-controlled tools are not the only avenue for scrutiny.

### v3

@@ 4.2. Standards Identification & Management
Auditors also used tools to formulate principles and norms to guide their investigations. While HCI research has not traditionally included frameworks and guidelines as tools for AI, Standards Identification & Management was a key focus for participants and comprised the largest collection of tools in our dataset. Standards Identification & Management ($N=206$), includes tools for Goal Articulation (e.g., broad principles statements), Self-Assessment (more specific procedural assessment tools, such as Microsoft’s AI Fairness Checklist (madaio_co-designing_2020)), Documentation (e.g., Model Cards (mitchell_model_2019)), Regulatory Awareness (tools, often paid services, for discovering and monitoring relevant regulations), Methods Design (standards for audit methodology), and Participatory Standard-Setting (methods for developing standards in collaboration with affected groups, such as Microsoft’s Community Jury (cass_community_2022)).
This category includes both internal organization standards and principles and formal national or international standards. Tools such as NIST’s Risk Management Framework (RMF) and relevant ISO standards are referenced extensively by auditing organizations (babl_auditor_certification), and regulators in Singapore and other nations have invested in similar tools as part of compliance and oversight frameworks (infocomm_media_development_authority_cataloguing_2023; kaye_risky_2023). While the weight and enforceability of these standards differ, they are united by a shared goal of defining audit methodologies and expectations for system performance.
@@
Need for more context-specific standards. Despite the large number of standards and evaluation frameworks surfaced in our tool survey, auditors still felt that evaluation frameworks needed refinement. Auditors emphasized the importance of standardized evaluation frameworks that provide clarity and consistency: “I think standardization is a big [concern]…” (P23) Many wished to streamline the auditing process by offering predefined structures and templates for assessment, which are essential for conducting audits effectively and facilitating communication. Most commonly found were goal-articulating “principle statements” ($N=86$), self-assessment checklists ($N=49$), and similar documents, while methods for participatory standard-setting ($N=5$) were comparatively rare in our dataset. The Standards Identification tools we found were particularly general in their applications, compared to other categories of tools—the principle statements, checklists, and similar resources we found were usually developed without a specific kind of target system in mind (Fig. E.4). Some participants found these tools too broad to easily apply:
@@
I think what I’ve seen is that companies and institutions… really, really struggle to understand, ‘What should we even do when it comes to auditing or evaluating the use of machine learning in our organizations?’ And while a template or a checklist is not the right answer, a lot of them don’t even know where to start… And so, (it helps) when you have a tool that…has some built in frameworks. (P8)
@@
Need for more standards beyond fairness. Some participants also thought that assessments templates and checklists were focused too narrowly on fairness assessment. One civil society auditor said, “I would love some guidance on audits generally and tools that describe the non-fairness components of an audit…” (P6) Another wished for resources that covered criteria such as explainability, privacy, and transparency: “Currently we have to.. find open source tools and put them together ourselves, and you need to have expertise to know what to look for” (P21).
@@
Need for clear and consistent regulatory guidance. Despite some participants’ desire for official frameworks like NIST’s Risk Management Framework (RMF) (tabassi_ai_2023), multiple participants commented on the difficulty of harmonizing current or expected regulatory guidance with practical implementation. Regulatory guidance itself may function as a tool, providing a framework for compliance with emerging policies. One civil society auditor asked:
@@
What are we evaluating for? And the question of, even when we have some kind of legal or other benchmark in mind, what are the metrics, and what are the benchmarks and other ways in which to evaluate, technical and otherwise, which also remain quite unclear? We’re seeing the ready adoption of audit language into policies, so that just kind of makes us nervous… What are we auditing for? (P2)
@@
Several emphasized the necessity of regulatory entities being more forthcoming in defining best practices. Some sought the guidance of the regulatory bodies in the domains they operate to establish their own frameworks but often struggle to translate industry expectations into meaningful standards for evaluation. Multiple participants felt there was a “culture or communications gap between the legal and compliance people on one side and the engineers on the other” (P12). As a result, some hoped for more collaborative approaches instead of command-and-control prescriptions: “Half the time we reach out [to regulators], and there’s just no one to contact… it’s just a black hole… ” (P13).
@@
Implications. Standards for evaluating AI systems must be simultaneously holistic, context-specific, inclusive, and compatible with practice. Some of the tools we found made advances in one or more of these dimensions, but few accomplished all three. Microsoft’s AI Fairness checklist, for example, was co-designed with practitioners (madaio_co-designing_2020) in an effort to be more compatible with practical challenges, but many of the other standards frameworks we found did not obviously consult practitioners and fewer involved affected stakeholders. Likewise, while NIST’s AI RMF includes safety, security, reliability, transparency, explainability, and privacy, in addition to fairness, its guidance for specific evaluation techniques remains fairly broad (tabassi_ai_2023).
@@
Research could continue to explore how regulatory standards could be translated into concrete metrics and other effective guidance for industry (wachter_bias_2021; guha_ai_2023). One participant at a large tech company, for example, preferred a “must, could, should” (P14) structure for regulatory guidance: a non-technical legal minimum (“must”) accompanied by more precise technical paths to compliance (“should”) and a set of ideal best practices for high performers and innovators (“could”).
@@ 4.3. Data Collection & Transparency Infrastructure
Gathering empirical evidence is a key step in AI auditing, but often poses the most significant challenge in practice. When model operators were unwilling or unable to release relevant documentation and other evidence, auditors turned to two main classes of tools to help.
@@
Tools for Transparency Infrastructure ($N=12$) are interfaces and databases hosted by model operators that allow controlled access to relevant data. This category includes tools for Structured or Application Programming Interface (API) Access (tools that allow auditors to interact with models and live systems, such as Google’s AI Test Kitchen (warkentin_join_2022)), tools for Data Sharing (platforms or trusts for hosting models and related data, such as the Gig Economy Data Hub (noauthor_gig_2021)), and tools for Secure & Private sharing (tools that help mitigate concerns with sharing data, such as Airbnb’s Project Lighthouse (airbnb_new_2020)).
@@
More commonly, though, tools for Data Collection ($N=47$), helped external auditors in particular gather information outside auditee-controlled interfaces. These tools help auditors gather data about model behavior, including relevant information not routinely collected by model operators. This category includes tools for Field Data Collection, which collect data from real systems and real users—including tools for Data Donation (such as Mozilla’s YouTube Regrets project (mozilla_foundation_youtube_2021)), Data Scraping (e.g., Tracking Exposed (agosti_tracking_2023)), Interviews/Surveys, and Compelled Transparency (e.g. tools such as MuckRock, which facilitates public records requests). We also found tools for Simulation (e.g. Meta’s Web Enabled Simulation platform for simulating interactions on Facebook (ahlgren_wes_2020)) and Bot Deployment (tools used for sock puppet auditing (bandy_problematic_2021), such as Selenium or Appium), both used to test systems with artificial or semi-artificial interactions.
Need for uncompromised data access. Data collection tools aim to address one of the challenges most frequently mentioned by our participants: the difficulty of accessing data and other vital information required to conduct meaningfully independent audits. Transparency Infrastructure tools provide external auditors with controlled access to models and data—especially for online platforms, in our dataset (Fig. E.4)—but they require investment from model operators. Despite court orders and regulations like Article 40 of the Digital Service Act (leerssen_digital_2023) that require the construction of transparency tools, one participant noted that key APIs used for auditing are becoming more costly and undependable:
There’s a direct impact on shutting off the access to information that affects people doing audits… we saw that with Reddit charging for their API and shutting down… Twitter charging astronomical, now, amounts for their API. It’s because everybody is scraping public Reddit and public Twitter to train large [AI] models… (P24)
They wished for “access to platform data… a context under which people can do controlled experiments, using data that is provided by platforms directly” (P24).
Auditors also said that corporate control over APIs undermined their independence in conducting audits which aligns with our definition of an audit (Def. B.2). A civil society auditor said,
I wish I had something to force people to give me their data… Part of the problem of auditing in general is that the only people who get let in are usually the people who are willing to say nice things about whatever the technology is being audited. (P6)
In practice, participants reported that key details, such as data sampling methods, data provenance, model versioning, metrics, and design justifications, were often omitted or only partially disclosed. And currently, the vetting process for API use often requires the auditor to disclose their intent for the evaluation in advance, which may compromise the integrity of the study.
I’m very much concerned that what’s going to happen is.. [platforms have] given all this access, and there will actually be… more of a cover up than there is now… They have so much power in this conversation to just share whatever information they want. (P7)
@@
Instead, one participant’s ideal was an “inspectability API” (P7), a required, standardized interface to allow researchers to interact with online platforms, including the ability to test different profiles, geographies, and other variables needed to evaluate disparate treatment, misinformation, and other algorithmic harms. Similarly, an auditor at a startup wished for centralized data archive available to the public:
@@
Challenges with independent data collection. Rather than rely on the model operator to provide access, auditors—especially external auditors—often turned to tools for Data Collection to obtain evidence themselves, sometimes developing and sharing their own tools and processes. Unlike the tools we found for Transparency Infrastructure, which were mostly not open source (25.0% open source in our dataset; see Fig. 3), tools for Data Collection were much more likely to be available under an open source license (68.1% open source in our dataset). Tools for Data Collection—most not built specifically for AI auditing—also comprised the most popular Github repositories in our dataset (Table E.2). For example, auditors used Selenium (noauthor_selenium_2023), a popular collection of open source tools for browser automation, to simulate user profiles while scraping data (known as a “sock puppet” audit (bandy_problematic_2021)).
(a) Number of open source tools in each taxonomy stage. Tools may be used in multiple stages.
[…]
@@
[…]
@@
Figure 3. Tool licensing by taxonomy stage (top) and by organization type (bottom).
While this approach gave auditors more freedom, it could also take more effort. Some of the tools we found—such as the Markup’s Citizen Browser (the_markup_citizen_2022), a data donation platform—were built from scratch to collect specific kinds of data for auditing. Existing tools for data scraping were helpful but often required extensive adaptation:
We almost always have to build custom scrapers to collect data… There’s some templates right out there for these scrapers, and then you usually have to customize them. And then there’s a huge amount of work to keep them alive… They break all the time with all these edge cases. And so they’re really a pain. I don’t really know that there’s a way to solve that. (P7)
Despite these difficulties, external auditors—particularly the journalists in our sample—still saw advantages in independent data collection. The same journalist noted:
@@
Challenges with understanding and processing data. Even when they had access to the data they wanted, participants noted that basic challenges involved with managing and analyzing data required more labor than any other task. A government auditor said, “90% of the work is figuring out what different tables are, and what different columns are, and working out whether it makes sense to join certain things. And then figuring out some meaningful metrics that we can draw from that data…” (P25). Participants also spent lots of time reviewing and requesting additional documents from audit targets.
Multiple participants wished for tools to help with tasks such as data collection and cleaning: “most of the value of data infrastructure is literally cleaning data” (P3). One auditor hoped for innovation in data quality management: “There are custom scrapers, a lot of human data quality work, and one thing that I have really wanted to do and never been able to do is try to figure out ways to get that data quality work done in more interesting ways” (P7).
@@
Risk of retaliation. Some external auditors also worried that external data collection tools—particularly tools for data scraping or data donation—may violate terms of service set by platforms and result in legal liability or retaliation. Auditors expressed concerns about legal risks from existing laws and regulations, particularly the Computer Fraud and Abuse Act (CFAA) and other privacy laws. Under the CFAA, for example, an auditor who violates a platforms’ terms of service may be held criminally liable.
It is difficult for auditors to determine what methods—such as public data scraping or even data donation—may be deemed “unauthorized” and criminal under the CFAA unless the audited organization grants authorization. In Sandvig v. Barr (2019), for example, the ACLU sued to allow researchers to set up false accounts (sock puppets) to audit computer algorithms (noauthor_sandvig_2020; american_civil_liberties_union_sandvig_2019). In 2021, Facebook used exactly this term of art—“unauthorized”—several times in its justification for disabling the accounts of a group of researchers auditing its advertising algorithms with a data donation tool called Ad Observer (clark_research_2021; edelson_we_2021). Facebook initially insinuated that disabling the researchers’ accounts was required by an Federal Trade Commission (FTC) consent decree, backtracking only after the FTC called the statement inaccurate (levine_letter_2021). One civil society auditor said,
Even if the legal risks have not been acted upon as much, there’s cases everyone points to in terms of attacks against researchers. It’s a matter of time before it ramps up. As soon as our work becomes threatening enough, that’s when it all really starts. (P24)
As a result, auditors engaged in external data collection had to take great steps to guard against liability and retaliation. One journalist said, “[The CFAA] is an incredible legal hangover for the type of work I do… how much lawyering I need to even get one tool off the ground is insane” (P7). Auditors also expressed hesitation about reforms that give platforms more control over what data is released. The same journalist continued,
@@
Implications. Future research could explore tools and processes for not only facilitating access to data—especially independently, through scraping or simulation—but also for ensuring data quality and integrity. Participants specifically wished for more tooling for data donation and user-driven auditing, a nascent area of research in human-computer interaction (lam_sociotechnical_2023; devos_toward_2022; deng_understanding_2023). Auditors also faced challenges common to data work in general, and research on practices surrounding data quality and data integrity may be applied specifically to discrimination testing and AI auditing methods.
Other challenges were more particular to auditing work. Future work could explore how auditors request information and interact with model operators. Transparency Infrastructure in particular is a nascent area of tooling that may become more common in auditing practice as AI regulation develops and as barriers to external data collection mount. Independent research may help guide these tools into more trustworthy mechanisms for disclosure, even as policymakers can ensure platform-controlled tools are not the only avenue for scrutiny.